Assignment 3(Chapter 3) To be submitted due to Dec.3 1. Each time a component is tested, the trial is a success (S) or failure (F). Suppose the component is tested repeatedly until a success occurs on three consecutive trials. Let Y denote the number of trials necessary to achieve this. List all outcomes corresponding to the five smallest possible values of Y, and state which Y value is associated with each one.
3.A mail-order computer business has six telephone lines. Let X denote the number of lines in use at a specified time Suppose the pmf of X is as given in the accompanying table. 0.10+0.15+0.20+0.25ー1020-t0.06-t604- Calculate the probability of each of the following events. a. fat most three lines are in use) b. (fewer than three lines are in use) c, {at least three lines are in use) d. (between two and five lines, inclusive, are in use) e. (between two and four lines, inclusive, are not in use) r. fat least four lines are not in use)
Answer #1

3. Here X denotes the number of lines in use at a specified time. Therefore, the probability P(X = x) = p(x).

a. P(at most three lines are in use)

0.10.15 + 0.20.25 0.7

b. P(fewer than three lines are in use)

= P(X < 3) p(0) + p(1) + p(2) 0.1 + 0.15 + 0.2 = 0.45

c. P(at least three lines are in use)

= 0.25 + 0.2 + 0.06 + 0.04 0.55

d. P(between two and five lines, inclusive, are in use)

-p(2) +p(3) +p(4) +p(5) 0.2+0.25+0.2 +0.06 0.71

e. P(between two and four lines, inclusive, are not in use) = 1 - P(between two and four lines, inclusive, are in use)

1-p(2) p(3) p(4)] 1 [0.2+0.250.2 = 1-0.65 0.35

f. P(at least four lines are not in use) = 1 - P(at least four lines are in use)

1-p(4) p(5) p(6)] 1-[0.2 0.060.04 = 1-0.3 0.70